# Unsupervised Domain Adaptation for Binary Classification with an Unobservable Source Subpopulation

### v2

@@ 7 Experiments
In this section, we study the Waterbirds dataset (Sagawa et al., 2019), which consists of 11,788 images. This public dataset is widely used to investigate spurious correlations in image classification. It is also well aligned with our problem setting of unsupervised domain adaptation under structured missingness, where specific combinations of labels and backgrounds are systematically absent in the labeled source domain, while labels are entirely unobserved in the target domain. The label $Y=1$ denotes a waterbird and $Y=0$ landbird. The background $A=1$ corresponds to a water background and $A=0$ a land background. It yields four label–background subpopulations, as summarized in Table 3.
@@
To construct a structured domain adaptation problem, we partition the full dataset into a source domain ($R=1$) and a target domain ($R=0$). Specifically, we allocate samples from three subgroups—$(Y=0,A=1)$, $(Y=1,A=0)$, and (Y=0,A=0)(Y=0,A=0)—into the source domain, with allocation rates denoted by parameters $a$, $b$, and $c$, respectively. The remaining subgroup, $(Y=1,A=1)$, is deliberately excluded from the source domain and appears only in the target domain. This setting reflects real-world scenarios in which a specific combination of label and background is structurally missing from labeled datasets due to systematic data collection biases or constraints. In the target domain, all four subgroups are retained, but the label variable $Y$ is treated as unobserved.
@@
To implement the proposed method, we apply the distribution matching approach to estimate the subclass proportions in the target domain. For feature extraction, we embed each image into a 512-dimensional feature vector using a ResNet-18 model (He et al., 2016) pre-trained on ImageNet (Deng et al., 2009), without additional fine-tuning. These embeddings serve as covariates ${\bf X}\in{\bf R}^{512}$ in our downstream analysis. Based on these feature vectors, we fit logistic regression models with $L_{2}$-regularization to estimate five key conditional probabilities required by both our proposed method and benchmark procedures: $\xi_{0}({\bf x})$, $\xi({\bf x})$, $\tau_{0}({\bf x})$, $\tau_{1}({\bf x})$ and $\kappa({\bf x})$.
@@
Figure 3: Performance comparison of our proposed estimators $\eta_{1}({\bf x})$, $\eta({\bf x})$, and the benchmark method $\xi_{1}({\bf x})$, $\xi({\bf x})$ under the setting $a=0.5$ with either $c=0.5$ and varying $b$ or $b=0.5$ and varying $c$.
@@
For empirical evaluation, we fix the subclass sampling rate at $a=0.5$ in the source domain and systematically vary the remaining subclass inclusion rates by setting either $b=0.5$ with $c\in\{0.1,0.2,\ldots,0.9\}$, or $c=0.5$ with $b\in\{0.1,0.2,\ldots,0.9\}$. For each configuration, the data generation process is repeated 50 times to account for sampling variability. We assess performance using two widely adopted classification metrics: accuracy and F1 score. Figure 3 presents boxplots summarizing the distribution of these metrics across repetitions for our proposed estimators $\widehat{\eta}_{1}({\bf x})$ and $\widehat{\eta}({\bf x})$, alongside benchmark estimators $\widehat{\xi}_{1}({\bf x})$ and $\widehat{\xi}({\bf x})$. The results demonstrate that the proposed methods consistently outperform the benchmarks in both accuracy and F1 score, suggesting enhanced robustness to structured subpopulation missingness in the unsupervised domain adaptation setting. Additional results are provided in Appendix S.3.
@@ 8 Discussions and Conclusions
SUPPLEMENT
S.1 Proofs and More Details in Section 4
@@
$\beta_{10}p_{10}({\bf x})+\beta_{00}p_{00}({\bf x})$ $=$ $\widetilde{\beta}_{10}p_{10}({\bf x})+\widetilde{\beta}_{00}p_{00}({\bf x}).$ (S.1)
@@
Since $\beta_{11}>0$, we obtain $p_{11}({\bf x})=\widetilde{p}_{11}({\bf x})$. Finally, (S.1) leads to $(\beta_{10}-\widetilde{\beta}_{10})p_{10}({\bf x})=(\widetilde{\beta}_{00}-\beta_{00})p_{00}({\bf x})$, which can only hold if $\beta_{10}=\widetilde{\beta}_{10}$ and $\widetilde{\beta}_{00}=\beta_{00}$ since $p_{10}({\bf x})\neq p_{00}({\bf x})$. This completes the proof. ∎
@@
$=$ $\int p({\bf x}|R=0,A=0){\rm log}\frac{p({\bf x}|R=0,A=0)}{\sum_{k=0}^{1}p({\bf x}|Y=k,A=0)\beta_{k0}\frac{1-\pi}{\hbox{pr}(R=0,A=0)}}{{\bf d}}{\bf x}$
$=$ $\int p({\bf x}|R=0,A=0){\rm log}\frac{p({\bf x}|R=0,A=0)}{\sum_{k=0}^{1}p({\bf x}|Y=k,A=0)\beta_{k0}\frac{1-\pi}{\hbox{pr}(R=0,A=0)}}{{\bf d}}{\bf x}$
$=$ $\int p({\bf x}|R=0,A=0){\rm log}\frac{p({\bf x}|R=0,A=0)}{p({\bf x}|R=1,A=0)}{{\bf d}}{\bf x}$
$-\int p({\bf x}|R=0,A=0){\rm log}\frac{\sum_{k=0}^{1}p({\bf x}|Y=k,A=0)\beta_{k0}\frac{1-\pi}{\hbox{pr}(R=0,A=0)}}{p({\bf x}|R=1,A=0)}{{\bf d}}{\bf x}$
$=$ $\int p({\bf x}|R=0,A=0){\rm log}\frac{p({\bf x}|R=0,A=0)}{p({\bf x}|R=1,A=0)}{{\bf d}}{\bf x}$
$-\int p({\bf x}|R=0,A=0){\rm log}\sum_{k=0}^{1}\frac{\hbox{pr}(Y=k|{\bf x},R=1,A=0)\beta_{k0}(1-\pi)\hbox{pr}(R=1,A=0)}{\hbox{pr}(R=1,Y=k,A=0)\hbox{pr}(R=0,A=0)}{{\bf d}}{\bf x}.$
@@
$\mbox{argmax}_{{\boldsymbol{\beta}}}E\left\{{\rm log}\sum_{k=0}^{1}\hbox{pr}(Y=k|{\bf x},R=1,A=0)\frac{\beta_{k0}}{\hbox{pr}(Y=k|R=1,A=0)}\middle|R=0,A=0\right\},$
@@
An Alternative Approach for Estimating $\boldsymbol{\beta}$}
In the main text, we explore the use of distribution matching for estimating $\boldsymbol{\beta}$}. Alternatively, it is sufficient to only consider some moments instead of the whole distribution. For any measurable function ${\bf m}({\bf x})$, the law of total expectation yields the identity:
$E\{{\bf m}({\bf x})\mid R=0,A=0\}\hbox{pr}(R=0,A=0)$ (S.2)
@@
Rewriting equation (S.2), we obtain the following linear system:
$(1-\pi)p_{0\cdot 0}^{-1}\left[E\{{\bf m}({\bf x})\mid 1,0\},E\{{\bf m}({\bf x})\mid 0,0\}\right]{\boldsymbol{\beta}}=E\{{\bf m}({\bf x})\mid R=0,A=0\},$
@@
${\boldsymbol{\beta}}=(1-\pi)^{-1}p_{0\cdot 0}\left[E\{{\bf m}({\bf x})\mid 1,0\},E\{{\bf m}({\bf x})\mid 0,0\}\right]^{-1}E\{{\bf m}({\bf x})\mid R=0,A=0\},$
provided that the $2\times 2$ matrix $\left[E\{{\bf m}({\bf x})\mid 1,0\},E\{{\bf m}({\bf x})\mid 0,0\}\right]$ is invertible. To use the idea of moment matching, one has the flexibility of choosing different moments ${\bf m}({\bf x})$. Certainly, a further research question of interest is to identify the optimal choice of this moment function, say, ${\bf m}_{\rm opt}({\bf x})$, by borrowing the semiparametric techniques (Bickel et al. 1993, Tsiatis 2006).
S.2 Proofs and More Details in Section 5
We define the Rademacher complexity (Bartlett and Mendelson 2002) that has been frequently used in machine learning literature to establish a generalization bound. Instead of considering the Rademacher complexity on $\mathcal{F}$ we define the class of weighted losses $\mathcal{G}(\ell,\mathcal{F})=[w(x,y)\ell\{g(x),y\}:g\in\mathcal{F}]$ and $n\in N$ we define its Rademacher complexity measure as
@@
For a probabilistic classifier: $\{\xi_{0}({\bf x}),1-\xi_{0}({\bf x})\}:\mathcal{X}\rightarrow\Delta^{2}$, and the parameter ${\boldsymbol{\beta}}^{\rm T}=(\beta_{10},\beta_{00})$ and $b_{1}=\hbox{pr}(Y=1|R=1,A=0)$, we define the centered logit function ${\boldsymbol{f}}:\mathcal{X}\rightarrow{\bf R}^{2}$ as $f_{0}({\bf x})={\rm log}\xi_{0}({\bf x})-\frac{1}{2}[{\rm log}\xi_{0}({\bf x})+{\rm log}\{1-\xi_{0}({\bf x})\}]$ and $f_{1}({\bf x})={\rm log}\{1-\xi_{0}({\bf x})\}-\frac{1}{2}[{\rm log}\xi_{0}({\bf x})+{\rm log}\{1-\xi_{0}({\bf x})\}]$. We define the functions $\mu(f_{0},b_{1})=\xi_{0}({\bf x})b_{1}^{-1}-\{1-\xi_{0}({\bf x})\}(1-b_{1})^{-1}$ and $\omega(f_{0},b_{1},\beta_{10},\varrho)=\xi_{0}({\bf x})b_{1}^{-1}\beta_{10}+\{1-\xi_{0}({\bf x})\}(1-b_{1})^{-1}(\varrho-\beta_{10})$, and notice that the objective is
@@
We decompose (S.2) using the Taylor expansion and obtain:
@@
It follows from Assumption 1 with probability at least $1-\delta$ it holds $\sup_{i\in[n_{0\cdot 0}]}\|\widehat{\boldsymbol{f}}({\bf x}_{i})-{\boldsymbol{f}}({\bf x}_{i})\|_{2}\leq cr_{n_{1\cdot 0}}\sqrt{{\rm log}(n_{0\cdot 0}){\rm log}(1/\delta)}$, we conclude that
@@
We combine it with the first order optimality condition for $\boldsymbol{\beta}$} to obtain
@@
$-(\widehat{\beta}_{10}-\beta_{10})\left\{\partial_{\beta_{10}}L(f_{0},b_{1},\widehat{\beta}_{10},\varrho)-\partial_{\beta_{10}}L(f_{0},b_{1},\beta_{10},\varrho)\right\}\geq\mu(\widehat{\beta}_{10}-\beta_{10})^{2}.$ (S.5)
@@
We notice that the event $\mathcal{E}$ has probability $1-5\delta$. Under the event there exists a $c_{8}>0$ such that the right-hand side in (S.2) is upper bounded as
@@
Combining the bounds (S.5) and (S.2) for left and right hand sides we obtain a $c_{10}>0$ such that on the event $\mathcal{E}$ it holds
@@
Further, sine $A_{i}$ is bound random variable, then we can obtain a constant $c_{9}>0$ such that for any $\delta>0$ with probability at least $1-\delta$ it holds
@@
$\|\widehat{\boldsymbol{\beta}}-{\boldsymbol{\beta}}\|_{1}\leq c_{10}\left\{r_{n_{1\cdot 0}}\sqrt{{\rm log}(n_{0\cdot 0}){\rm log}(1/\delta)}+\sqrt{\frac{{\rm log}(1/\delta)}{n_{1\cdot 0}}}+\sqrt{\frac{{\rm log}(1/\delta)}{n_{0}}}++\sqrt{\frac{{\rm log}(1/\delta)}{n_{0\cdot 0}}}\right\}.$
@@
(Derivatives).Th following holds:
@@
$=$ ${\mathrm{E}}\left[\ell\{h({\bf X}),Y\}w(Y)\big|A=0,R=1\right]=:{\mathcal{L}}_{1}(h,w).$
@@
Uniform bound on (a) To control (a) in (S.2) we establish a concentration bound on the following generalization error
@@
$F({\bf Z}_{1:n_{1\cdot 0}})\leq E\{F({\bf Z}_{1:n_{1\cdot 0}})\}+c_{1}\sqrt{\frac{{\rm log}(1/\delta)}{n_{1\cdot 0}}}.$ (S.8)
@@
$E\{F({\bf Z}_{1:n_{1\cdot 0}})\}\leq 2\mathcal{R}_{n_{1\cdot 0}}(\mathcal{G}).$ (S.9)
Combining (S.8) and (S.9) we obtain
$(a)\leq 2\mathcal{R}_{n_{1\cdot 0}}(\mathcal{G})+c_{1}\sqrt{\frac{{\rm log}(1/\delta)}{n_{1\cdot 0}}}$ (S.10)
@@
$\leq$ $\|\ell_{g}\|_{\infty}\left\{C_{1}\|\widehat{\boldsymbol{\beta}}-{\boldsymbol{\beta}}\|_{1}+C_{2}\left(\sqrt{\frac{{\rm log}(1/\delta)}{n_{0}}}+\sqrt{\frac{{\rm log}(1/\delta)}{n_{1}}}\right)+c_{2}\sqrt{\frac{{\rm log}(1/\delta)}{n_{1\cdot 0}}}\right\}$
@@
$\widehat{\mathcal{L}}_{1}(h,w)-{\mathcal{L}}_{1}(h,w)\leq c_{3}\sqrt{\frac{{\rm log}(1/\delta)}{n_{1\cdot 0}}}.$ (S.12)
Finally, using (S.10) on (a) (which is true on an event of probability $\geq 1-\delta$), (S.2) on (b) and (c) (simultaneously true on an event of probability $1-5\delta$), and (S.12) on (d) (holds on an event of probability $\geq 1-\delta$) we conclude that with probability at least $1-7\delta$ the following holds
${\mathcal{L}}_{0}(\widehat{h}_{\widehat{w}})-{\mathcal{L}}_{0}(h)\leq 2\mathcal{R}_{n_{1\cdot 0}}(\mathcal{G})+CB\|\widehat{\boldsymbol{\beta}}-{\boldsymbol{\beta}}\|_{1}+c\left\{\sqrt{\frac{{\rm log}(1/\delta)}{n_{1\cdot 0}}}+\sqrt{\frac{{\rm log}(1/\delta)}{n_{0}}}+\sqrt{\frac{{\rm log}(1/\delta)}{n_{1}}}\right\}.$
@@
$|\widehat{w}(y)-w(y)|\leq CB_{1}\|\widehat{\boldsymbol{\beta}}-{\boldsymbol{\beta}}\|_{1}+CB_{1}(c_{1}+c_{2})\left(\sqrt{\frac{{\rm log}(1/\delta)}{n_{0}}}+\sqrt{\frac{{\rm log}(1/\delta)}{n_{1}}}\right).$
@@
$\leq$ $CB_{1}(\|\widehat{\boldsymbol{\beta}}-{\boldsymbol{\beta}}\|_{1}+|\widehat{\alpha}_{00}-\alpha_{00}|+|\widehat{\alpha}_{10}-\alpha_{10}|).$

### v3

@@ 7 Experiments
In this section, we analyze the Waterbirds dataset (Sagawa et al., 2019), which consists of 11,788 images. This public dataset is widely used to investigate spurious correlations in image classification. It is also well aligned with our problem setting of unsupervised domain adaptation under structured missingness, where specific combinations of labels and backgrounds are systematically absent in the labeled source domain, while labels are entirely unobserved in the target domain. The label $Y=1$ denotes a waterbird and $Y=0$ landbird. The background $A=1$ corresponds to a water background and $A=0$ a land background. It yields four label–background subpopulations, as summarized in Table 3.
@@
To construct a structured domain adaptation problem, we partition the full dataset into a source domain ($R=1$) and a target domain ($R=0$). Specifically, we allocate samples from three subgroups, $(Y=0,A=1)$, $(Y=1,A=0)$ and $(Y=0,A=0)$, into the source domain, with allocation rates denoted by parameters $a$, $b$, and $c$, respectively. The remaining subgroup, $(Y=1,A=1)$, is deliberately excluded from the source domain and appears only in the target domain. This setting reflects real-world scenarios in which a specific combination of label and background is structurally missing from labeled datasets due to systematic data collection biases or constraints. In the target domain, all four subgroups are retained, but the label variable $Y$ is treated as unobserved.
@@
To implement the proposed method, we apply the distribution matching approach to estimate the subclass proportions in the target domain. For feature extraction, we embed each image into a 512-dimensional feature vector using a ResNet-18 model (He et al., 2016) and a ViT-16 model (Heo et al., 2021), both pre-trained on ImageNet (Deng et al., 2009), without additional fine-tuning. These embeddings serve as covariate ${\bf X}\in{\bf R}^{512}$ in our downstream analysis. Based on these feature vectors, we fit logistic regression models with $L_{2}$-regularization to estimate five key conditional probabilities required by both our proposed method and benchmark procedures: $\xi_{0}({\bf x})$, $\xi({\bf x})$, $\tau_{0}({\bf x})$, $\tau_{1}({\bf x})$ and $\kappa({\bf x})$.
@@
Figure 2: Performance comparison of our proposed estimator $\eta({\bf x})$, and the naive methods $\xi({\bf x})$ and $\gamma({\bf x})$ under the setting $a=0.7$ with either $c=0.5$ and varying $b$ or $b=0.5$ and varying $c$.
Figure 3: Performance comparison of our proposed estimator $\eta({\bf x})$, and the naive methods $\xi({\bf x})$ and $\gamma({\bf x})$ under the setting $a=0.5$ with either $c=0.5$ and varying $b$ or $b=0.5$ and varying $c$.
Figure 4: Performance comparison of our proposed estimator $\eta({\bf x})$, and the naive methods $\xi({\bf x})$ and $\gamma({\bf x})$ under the setting $a=0.3$ with either $c=0.5$ and varying $b$ or $b=0.5$ and varying $c$.
@@
For empirical evaluation, we fix the subclass sampling rate at $a=0.7$ in the source domain and systematically vary the remaining subclass inclusion rates by setting either $b=0.5$ with $c\in\{0.1,0.2,\ldots,0.9\}$, or $c=0.5$ with $b\in\{0.1,0.2,\ldots,0.9\}$. For each configuration, the data generation process is repeated 50 times to account for sampling variability. We assess performance using two widely adopted classification metrics: accuracy and F1 score. Figure 2 presents boxplots summarizing the distributions of the metrics across repeated experiments for our proposed estimator $\widehat{\eta}({\bf x})$ and the two naive estimators $\widehat{\xi}({\bf x})$ and $\widehat{\gamma}({\bf x})$. The left panel shows the results based on features extracted by ResNet-18, while the right panel shows the results based on features extracted by ViT-16. We also reduce the proportion $a$ from $0.7$ to $0.5$ and $0.3$, and similarly demonstrate the results in Figure 3 and Figure 4, respectively.
The observations and conclusions from this experiment can be summarized as follows. First, the proposed correct predictive probability $\eta({\bf x})$ consistently achieves the best performance across most scenarios. This is particularly evident when the subpopulation $(Y=0,A=1)$ constitutes a larger proportion $(a=0.7)$ in the source domain (Figure 2). This behavior is logical; because $(Y=0,A=1)$ is the only class containing information about $A=1$, a higher representation of this subgroup helps compensate for the absence of the $(Y=1,A=1)$ subgroup. Second, the performance of the proposed $\eta({\bf x})$ can degrade, even falling behind the two incorrect naive benchmarks, if the proportion of either $(Y=1,A=0)$ or $(Y=0,A=0)$ becomes exceedingly small either in the source domain or in the target. In such cases, the source domain may effectively suffer from multiple missing subgroups (beyond just $(Y=1,A=1)$), elevating the problem’s difficulty well beyond the scope of this paper. Finally, choosing ViT-16 generally brings much better performance than using ResNet-18, especially for the two naive benchmarks.
@@ 8 Discussions and Conclusions
The invariance assumption imposed in (3) can be viewed as a conditional, or more nuanced, extension of the standard label shift assumption. While standard label shift assumes that the distribution of $\bf X$} has no change across domains conditional on the label $Y$, equation (3) postulates that its distribution remains the same conditional on both the label $Y$ and the environment $A$. As discussed in Section 3, this assumption is justifiable in a variety of practical contexts. In practice, however, the environment $A$ itself may undergo shifts. For instance, the source domain might contain only two backgrounds (e.g., water and land), whereas the target domain introduces a third (e.g., sky). This scenario effectively transitions into an out-of-distribution (OOD) generalization problem, where the shift occurs with respect to the environment $A$ rather than the label $Y$. Addressing a missing subgroup in the source domain within this OOD context extends significantly beyond the scope of this paper, as the model identification challenges discussed in Section 4.3 become substantially more intractable. Consequently, this problem remains an important area for future work.
Broader Impact Statement
This work addresses a practical limitation of unsupervised domain adaptation by enabling reliable prediction when an entire subpopulation is absent from the source data, a situation that commonly arises in healthcare, ecological monitoring, and other domains where certain subgroups are systematically underrepresented due to data collection constraints. By explicitly modeling structured missingness, our approach can help reduce biased predictions on groups that conventional methods would otherwise misclassify, contributing to fairer and more robust machine learning systems.
Acknowledgments
Sharon Li is supported in part by the AFOSR Young Investigator Program under award number FA9550-23-1-0184, National Science Foundation (NSF) under awards IIS-2237037 and IIS-2331669, Office of Naval Research under grant number N00014-23-1-2643, Schmidt Sciences Foundation, Open Philanthropy, Alfred P. Sloan Fellowship, and gifts from Google and Amazon. Jiwei Zhao is supported in part by NSF under awards DMS-1953526, DMS-2122074 and DMS-2310942, and National Institutes of Health under award R01DC021431. Any opinions, findings, and conclusions or recommendations expressed in this paper are those of the author(s) and do not necessarily reflect the views of the funding agencies.
Appendix A Appendix
A.1 Proofs and More Details in Section 4
@@
$\beta_{10}p_{10}({\bf x})+\beta_{00}p_{00}({\bf x})$ $=$ $\widetilde{\beta}_{10}p_{10}({\bf x})+\widetilde{\beta}_{00}p_{00}({\bf x}).$ (14)
@@
Since $\beta_{11}>0$, we obtain $p_{11}({\bf x})=\widetilde{p}_{11}({\bf x})$. Finally, (14) leads to $(\beta_{10}-\widetilde{\beta}_{10})p_{10}({\bf x})=(\widetilde{\beta}_{00}-\beta_{00})p_{00}({\bf x})$, which can only hold if $\beta_{10}=\widetilde{\beta}_{10}$ and $\widetilde{\beta}_{00}=\beta_{00}$ since $p_{10}({\bf x})\neq p_{00}({\bf x})$. This completes the proof. ∎
@@
$=$ $\int p({\bf x}|R=0,A=0){\rm log}\frac{p({\bf x}|R=0,A=0)}{\sum_{k=0}^{1}p({\bf x}|Y=k,A=0)\beta_{k0}\frac{1-\pi}{\hbox{pr}(R=0,A=0)}}\mathrm{d}{\bf x}$
$=$ $\int p({\bf x}|R=0,A=0){\rm log}\frac{p({\bf x}|R=0,A=0)}{p({\bf x}|R=1,A=0)}\mathrm{d}{\bf x}$
$-\int p({\bf x}|R=0,A=0){\rm log}\frac{\sum_{k=0}^{1}p({\bf x}|Y=k,A=0)\beta_{k0}\frac{1-\pi}{\hbox{pr}(R=0,A=0)}}{p({\bf x}|R=1,A=0)}\mathrm{d}{\bf x}$
$=$ $\int p({\bf x}|R=0,A=0){\rm log}\frac{p({\bf x}|R=0,A=0)}{p({\bf x}|R=1,A=0)}\mathrm{d}{\bf x}$
$-\int p({\bf x}|R=0,A=0){\rm log}\sum_{k=0}^{1}\frac{\hbox{pr}(Y=k|{\bf x},R=1,A=0)\beta_{k0}(1-\pi)\hbox{pr}(R=1,A=0)}{\hbox{pr}(R=1,Y=k,A=0)\hbox{pr}(R=0,A=0)}\mathrm{d}{\bf x}.$
@@
$\mbox{argmax}_{{\bm{\beta}}}E\left\{{\rm log}\sum_{k=0}^{1}\hbox{pr}(Y=k|{\bf x},R=1,A=0)\frac{\beta_{k0}}{\hbox{pr}(Y=k|R=1,A=0)}\middle|R=0,A=0\right\},$
@@
An Alternative Approach for Estimating $\bm{\beta}$}
In the main text, we explore the use of distribution matching for estimating $\bm{\beta}$}. Alternatively, it is sufficient to only consider some moments instead of the whole distribution. For any measurable function ${\bf m}({\bf x})$, the law of total expectation yields the identity:
$E\{{\bf m}({\bf x})\mid R=0,A=0\}\hbox{pr}(R=0,A=0)$ (15)
@@
Rewriting equation (15), we obtain the following linear system:
$(1-\pi)p_{0\cdot 0}^{-1}\left[E\{{\bf m}({\bf x})\mid 1,0\},E\{{\bf m}({\bf x})\mid 0,0\}\right]{\bm{\beta}}=E\{{\bf m}({\bf x})\mid R=0,A=0\},$
@@
${\bm{\beta}}=(1-\pi)^{-1}p_{0\cdot 0}\left[E\{{\bf m}({\bf x})\mid 1,0\},E\{{\bf m}({\bf x})\mid 0,0\}\right]^{-1}E\{{\bf m}({\bf x})\mid R=0,A=0\},$
provided that the $2\times 2$ matrix $\left[E\{{\bf m}({\bf x})\mid 1,0\},E\{{\bf m}({\bf x})\mid 0,0\}\right]$ is invertible. To use the idea of moment matching, one has the flexibility of choosing different moments ${\bf m}({\bf x})$. Certainly, a further research question of interest is to identify the optimal choice of this moment function, say, ${\bf m}_{\rm opt}({\bf x})$, by borrowing the semiparametric techniques (Bickel et al., 1993; Tsiatis, 2006).
A.2 Proofs and More Details in Section 5
We define the Rademacher complexity (Bartlett and Mendelson, 2002) that has been frequently used in machine learning literature to establish a generalization bound. Instead of considering the Rademacher complexity on $\mathcal{F}$ we define the class of weighted losses $\mathcal{G}(\ell,\mathcal{F})=[w(x,y)\ell\{g(x),y\}:g\in\mathcal{F}]$ and $n\in N$ we define its Rademacher complexity measure as
@@ A.2 Proofs and More Details in Section 5
For a probabilistic classifier: $\{\xi_{0}({\bf x}),1-\xi_{0}({\bf x})\}:\mathcal{X}\rightarrow\Delta^{2}$, and the parameter ${\bm{\beta}}^{\rm T}=(\beta_{10},\beta_{00})$ and $b_{1}=\hbox{pr}(Y=1|R=1,A=0)$, we define the centered logit function ${\bm{f}}:\mathcal{X}\rightarrow{\bf R}^{2}$ as $f_{0}({\bf x})={\rm log}\xi_{0}({\bf x})-\frac{1}{2}[{\rm log}\xi_{0}({\bf x})+{\rm log}\{1-\xi_{0}({\bf x})\}]$ and $f_{1}({\bf x})={\rm log}\{1-\xi_{0}({\bf x})\}-\frac{1}{2}[{\rm log}\xi_{0}({\bf x})+{\rm log}\{1-\xi_{0}({\bf x})\}]$. We define the functions $\mu(f_{0},b_{1})=\xi_{0}({\bf x})b_{1}^{-1}-\{1-\xi_{0}({\bf x})\}(1-b_{1})^{-1}$ and $\omega(f_{0},b_{1},\beta_{10},\varrho)=\xi_{0}({\bf x})b_{1}^{-1}\beta_{10}+\{1-\xi_{0}({\bf x})\}(1-b_{1})^{-1}(\varrho-\beta_{10})$, and notice that the objective is
@@
We decompose (A.2) using the Taylor expansion and obtain:
@@
It follows from Assumption 1 with probability at least $1-\delta$ it holds $\sup_{i\in[n_{0\cdot 0}]}\|\widehat{\bm{f}}({\bf x}_{i})-{\bm{f}}({\bf x}_{i})\|_{2}\leq cr_{n_{1\cdot 0}}\sqrt{{\rm log}(n_{0\cdot 0}){\rm log}(1/\delta)}$, we conclude that
@@
We combine it with the first order optimality condition for $\bm{\beta}$} to obtain
@@
$-(\widehat{\beta}_{10}-\beta_{10})\left\{\partial_{\beta_{10}}L(f_{0},b_{1},\widehat{\beta}_{10},\varrho)-\partial_{\beta_{10}}L(f_{0},b_{1},\beta_{10},\varrho)\right\}\geq\mu(\widehat{\beta}_{10}-\beta_{10})^{2}.$ (18)
@@
We notice that the event $\mathcal{E}$ has probability $1-5\delta$. Under the event there exists a $c_{8}>0$ such that the right-hand side in (A.2) is upper bounded as
@@
Combining the bounds (18) and (A.2) for left and right hand sides we obtain a $c_{10}>0$ such that on the event $\mathcal{E}$ it holds
@@
Further, since $A_{i}$ is bound random variable, then we can obtain a constant $c_{9}>0$ such that for any $\delta>0$ with probability at least $1-\delta$ it holds
@@
$\|\widehat{\bm{\beta}}-{\bm{\beta}}\|_{1}\leq c_{10}\left\{r_{n_{1\cdot 0}}\sqrt{{\rm log}(n_{0\cdot 0}){\rm log}(1/\delta)}+\sqrt{\frac{{\rm log}(1/\delta)}{n_{1\cdot 0}}}+\sqrt{\frac{{\rm log}(1/\delta)}{n_{0}}}++\sqrt{\frac{{\rm log}(1/\delta)}{n_{0\cdot 0}}}\right\}.$
@@
(Derivatives).The following equations hold:
@@
$=$ ${\mathrm{E}}\left[\ell\{h({\bf X}),Y\}w(Y)\big|A=0,R=1\right]=:\L _{1}(h,w).$
@@
Uniform bound on (a) To control (a) in (A.2) we establish a concentration bound on the following generalization error
@@
$F({\bf Z}_{1:n_{1\cdot 0}})\leq E\{F({\bf Z}_{1:n_{1\cdot 0}})\}+c_{1}\sqrt{\frac{{\rm log}(1/\delta)}{n_{1\cdot 0}}}.$ (21)
@@
$E\{F({\bf Z}_{1:n_{1\cdot 0}})\}\leq 2\mathcal{R}_{n_{1\cdot 0}}(\mathcal{G}).$ (22)
Combining (21) and (22) we obtain
$(a)\leq 2\mathcal{R}_{n_{1\cdot 0}}(\mathcal{G})+c_{1}\sqrt{\frac{{\rm log}(1/\delta)}{n_{1\cdot 0}}}$ (23)
@@
$\leq$ $\|\ell_{g}\|_{\infty}\left\{C_{1}\|\widehat{\bm{\beta}}-{\bm{\beta}}\|_{1}+C_{2}\left(\sqrt{\frac{{\rm log}(1/\delta)}{n_{0}}}+\sqrt{\frac{{\rm log}(1/\delta)}{n_{1}}}\right)+c_{2}\sqrt{\frac{{\rm log}(1/\delta)}{n_{1\cdot 0}}}\right\}$
@@
$\widehat{\mathcal{L}}_{1}(h,w)-{\mathcal{L}}_{1}(h,w)\leq c_{3}\sqrt{\frac{{\rm log}(1/\delta)}{n_{1\cdot 0}}}.$ (25)
Finally, using (23) on (a) (which is true on an event of probability $\geq 1-\delta$), (A.2) on (b) and (c) (simultaneously true on an event of probability $1-5\delta$), and (25) on (d) (holds on an event of probability $\geq 1-\delta$) we conclude that with probability at least $1-7\delta$ the following holds
${\mathcal{L}}_{0}(\widehat{h}_{\widehat{w}})-{\mathcal{L}}_{0}(h)\leq 2\mathcal{R}_{n_{1\cdot 0}}(\mathcal{G})+CB\|\widehat{\bm{\beta}}-{\bm{\beta}}\|_{1}+c\left\{\sqrt{\frac{{\rm log}(1/\delta)}{n_{1\cdot 0}}}+\sqrt{\frac{{\rm log}(1/\delta)}{n_{0}}}+\sqrt{\frac{{\rm log}(1/\delta)}{n_{1}}}\right\}.$
@@
$|\widehat{w}(y)-w(y)|\leq CB_{1}\|\widehat{\bm{\beta}}-{\bm{\beta}}\|_{1}+CB_{1}(c_{1}+c_{2})\left(\sqrt{\frac{{\rm log}(1/\delta)}{n_{0}}}+\sqrt{\frac{{\rm log}(1/\delta)}{n_{1}}}\right).$
@@
$\leq$ $CB_{1}(\|\widehat{\bm{\beta}}-{\bm{\beta}}\|_{1}+|\widehat{\alpha}_{00}-\alpha_{00}|+|\widehat{\alpha}_{10}-\alpha_{10}|).$